OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4). A. y = four thirdsx + 8 B. y = four thirdsx + 3 C. y = -three fourthsx + 8 D. y = -three fourthsx + 3 @JoshR1213

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Get the original equation into y=mx+b form first

OpenStudy (nurali):

4x-3y= -12 -3y= -12 -4x y= 4 +4/3x In other words, your new equation is y= 4/3x +b So using y= 4/3x +b and the point (-3,4): 4= 4/3(-3/1) +B 4= -12/3 +b 4= -4 +b b= 8 Therefore y= 4/3x +8

OpenStudy (anonymous):

put in y=mx+b I always use this for equations like that